OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5). You must show all work.

OpenStudy (anonymous):

2x – 5 = 3(2x + 5) 2x -5 = 6x + 15 -15 -5 = 6x - 2x 4x = -20 ->x = -20/4 = -

OpenStudy (anonymous):

x = -5